1. A gas takes up a volume of 17 liters, has a pressure of 2.3 atm, and a temperature of 215 K. If I raise the temperature to 350 K and lower the pressure to 1.5 atm, what is the new volume of the gas?
2. I have an unknown volume of gas at a pressure of 0.5 atm and a temperature of 325 K. If I raise the pressure to 1.2 atm, decrease the temperature to 320 K, and measure the final volume to be 48 liters, what was the initial volume of the gas?
3. A gas at a temperature of 10°C and a volume of 4.5 liters has a pressure of 1.5 atm. If the temperature is raised to 29°C and the pressure goes to 2.8 atm, what is the new volume of the gas?
4. If I have 15 liters of gas held at a pressure of 68 atm and a temperature of 700 K, what will be the volume of the gas if I decrease the pressure to 35 atm and decrease the temperature to 450 K?
1
Expert's answer
2020-04-08T13:33:31-0400

In order to solve these problems I'll use the general gas equation:

p1V1/T1 = p2V2/T2;


  1. V2=V1*p1*T2/(p2*T1)=17*2.3*350/(1.5*215)=42.43 L
  2. V1=T1*p2*V2/(T2*p1)=325*1.2*48/(320*0.5)=117 L
  3. T1=273+10=283 K; T2=273+29=302K;

V2=V1*p1*T2/(p2*T1)=4.5*1.5*302/(2.8*283)=2.57 L

4. V2=V1*p1*T2/(p2*T1)= 15*68*450/(35*700)=18.73 L